### v1

@@ 3 Stackelberg value is unachievable
[…]
@@
It suffices to prove this lower bound in a centralized environment where a single learner can choose action pairs $(a,b)$ and observes rewards for both players (Lemma C.1). We show that on the instances $\mathcal{I}$ and $\tilde{\mathcal{I}}$ in Table 1 (with $\delta=O(1/\sqrt{T})$), at least one of the players incurs linear regret on at least one of the instances. The small value of $\delta$ means that the algorithm cannot distinguish between these instances with constant probability. Nonetheless, the benchmarks are very different: on instance $\mathcal{I}$, $(a^{*},b^{*})=(a_{1},b_{1})$, $\beta_{1}^{\text{orig}}=0.6$ and $\beta_{2}^{\text{orig}}=\delta>0$; on instance $\tilde{\mathcal{I}}$, $(a^{*},b^{*})=(a_{2},b_{1})$, $\beta_{1}^{\text{orig}}=0.5$, and $\beta_{2}^{\text{orig}}=0.6$. Intuitively, when the algorithm fails to distinguish between these instances, then it must choose the same distribution over $\mathcal{A}\times\mathcal{B}$ on both $\mathcal{I}$ and $\tilde{\mathcal{I}}$. However, any such distribution either incurs constant loss for the leader on $\mathcal{I}$ or constant loss for the follower on $\tilde{\mathcal{I}}$. We formalize this proof in Section C.4. ∎
@@ 4.1 $\gamma$-tolerant benchmark
Defining the $\epsilon$-best response set $\mathcal{A}_{\epsilon}$ for the leader is more subtle. Informally, we define this set to include any action $a\in\mathcal{A}$ which has “any chance” of doing at least as well as the the leader’s best action if the follower is $\epsilon$-best responding. Specifically, this includes actions $a$ where some action $b\in\mathcal{B}_{\epsilon}(a)$ achieves utility close to $\max_{a^{\prime}\in\mathcal{A}}\min_{b^{\prime}\in\mathcal{B}_{\epsilon}(a)}v_{1}(a^{\prime},b^{\prime})$ for the leader:
$\mathcal{A}_{\epsilon}=\{a\in\mathcal{A}\mid\max_{b\in\mathcal{B}_{\epsilon}(a)}v_{1}(a,b)\geq\max_{a^{\prime}\in\mathcal{A}}\min_{b^{\prime}\in\mathcal{B}_{\epsilon}(a^{\prime})}v_{1}(a^{\prime},b^{\prime})-\epsilon\}$
@@ 4.2 Linear regret for ExploreThenCommit
We first show that out-of-box stochastic bandit algorithms do not directly provide sublinear regret against the $\gamma$-tolerant benchmark, where the challenge is that the leader’s learning gets distorted when both players simultaneously learn. To demonstrate this, we consider ExploreThenCommit (Algorithm 1) and show that if both the leader and follower are running this algorithm in the strongly decentralized setting, the regret could be linear for both players.
@@
             Set $S(c):=\left\{r\mid\exists(t^{\prime},c_{t^{\prime}},r)\in H^{*}\text{ s.t. }c=c_{t^{\prime}}\right\}$   // observed rewards
             $\hat{v}(c)\leftarrow(\sum_{r\in S(c)}r)/|S(c)|$   // compute empirical mean
@@
[…]
@@
The intuition is that in the leader’s exploration phase, the follower alternates uniformly between all actions $\mathcal{B}$. This distorts the leader’s learning during the leader’s exploration phase, and as a result, the leader can choose a highly suboptimal arm during the commit phase. This can lead to linear regret for both players. We construct a single instance (Table 2, with $\delta=0.1$) where both players incur linear regret. The full proof is deferred to Section D.1. ∎
4.3 Warmup: Simple Pair of Algorithms Achieving Sublinear Regret
The constant regret in Proposition 4.3 motivates the design of more sophisticated algorithms where the leader waits for the follower to partially converge before starting to learn. As a warmup, we show that a simple modification of the setup of Proposition 4.3 guarantees sublinear regret (i.e., $O\left(|\mathcal{A}|^{1/3}|\mathcal{B}|^{1/3}(\log T)^{1/3}T^{2/3}\right)$ regret for both players. Both players run ExploreThenCommit-based algorithms, but the leader waits for the follower to finish exploring before starting to explore. More specifically, the leader runs ExploreThenCommitThrowOut, which acts similar to ExploreThenCommit, but with an extra exploration phase at the start, after which all rewards are thrown out. This initial phase is to allow the follower to partially converge.
@@ 4.3 Warmup: Simple Pair of Algorithms Achieving Sublinear Regret
       $H^{*}=\left\{(t^{\prime}-E^{\prime}\cdot|\mathcal{C}|,c_{t^{\prime}},r)\mid\exists(t^{\prime},c_{t^{\prime}},r)\in H\text{ s.t. }t^{\prime}>E^{\prime}\cdot|\mathcal{C}|\right\}$   // Throw out first $E^{\prime}\cdot|\mathcal{C}|$ rounds of history
@@
The “throw out” phase for $\texttt{ALG}_{1}$ enables $\texttt{ALG}_{2}$ to learn and commit to near-optimal actions. The meaningful exploration for $\texttt{ALG}_{1}$ thus begins after the follower has committed to actions. This enables $\texttt{ALG}_{1}$ to identify a near-optimal action given the arms that $\texttt{ALG}_{2}$ has already committed to after the first phase of exploration. Returning to our $\gamma$-tolerant benchmarks, for each player, we can upper bound regret by setting $\epsilon$ to be the suboptimality of the other player and achieve the desired regret bound. The full proof is deferred to Appendix D.2.
@@
4.4 Main Result: Richer Algorithms Achieving Sublinear Regret
@@
The algorithm $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$ operates in the strongly decentralized setting, taking as input $E\in[T/\left|\mathcal{A}\right|]$. When $\texttt{ALG}_{1}$ applied to an instance, for the first $|\mathcal{A}|\cdot E$ rounds, the algorithm $\texttt{ALG}_{1}$ pulls each arm in $\mathcal{A}$ a total of $E$ times, and then discards all history from these rounds. For the remaining $T-|\mathcal{A}|\cdot E$ rounds, the algorithm runs UCB, computing the upper confidence bound $v^{\text{UCB}}_{1}(a)=\hat{v}_{1,t}(a)+\alpha_{t}(a)$ using confidence bound $\alpha_{t}(a)=\Theta\left(\sqrt{(\log T)/n_{E\cdot|\mathcal{A}|,t}(a)}\right)$, where $\hat{v}_{1,t}(a)$ is the empirical mean and $n_{E\cdot|\mathcal{A}|,t}(a)$ is the number of times that action $a$ is chosen in the UCB phase after time step $E\cdot|\mathcal{A}|$ and prior to time step $t$. The algorithm then chooses the arm with maximum upper confidence bound.
@@
       $H^{*}=\left\{(t^{\prime}-E\cdot|\mathcal{A}|,a_{t^{\prime}},r)\mid\exists(t^{\prime},a_{t^{\prime}},r)\in H\text{ s.t. }t^{\prime}>E\cdot|\mathcal{A}|\right\}$   // Throw out first $E\cdot|\mathcal{A}|$ rounds of history
       Initialize $\hat{v}_{1}(a)=1$ for $a\in\mathcal{A}$.   // Initialize empirical means.
       Initialize $v^{\text{UCB}}_{1}(a)=1$ for $a\in\mathcal{A}$.   // Initialize UCB.
@@
             Set $S(a):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}}))\in H^{*}\text{ s.t. }a=a_{t^{\prime}},r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}})=r\right\}$   // Observed rewards
@@
                   $\hat{v}_{1}(a)\leftarrow(\sum_{r\in S(a)}r)/|S(a)|$   // Empirical mean
                   $\alpha(a)\leftarrow 10\cdot\frac{\sqrt{\log T}}{\sqrt{|S(a)|}}$   // confidence bound width
@@ 10                   $v^{\text{UCB}}_{1}(a)\leftarrow\min(1,\hat{v}_{1}(a)+\alpha(a))$
      Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\left(v^{\text{UCB}}_{1}(a)\right)$.   // arm with max upper confidence bound
@@
[…]
@@
The intuition is that the exploration phase of ExploreThenUCB ensures that all of the follower’s actions have bounded suboptimality, and the UCB phase accounts for the follower changing which action they choose over time. In more detail, high-probability instantaneous regret guarantees that after the explore phase, all actions that the follower’s chooses are within the $\epsilon$-best-response set $\mathcal{B}_{\epsilon^{*}}(a)$ for $\epsilon^{*}=\Theta((|\mathcal{A}|\cdot|\mathcal{B}|\cdot\log T)^{1/3}T^{-1/3})$. For the UCB phase, the main lemma (Lemma D.4) is that if an arm $a\in\mathcal{A}$ is pulled, the empirical mean is at least $\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\Theta\left(\sqrt{\log T/n_{E\cdot|\mathcal{A}|,t}(a)}\right)$ (the optimal utility for the leader when the follower worst-case $\epsilon$-best-responds minus the confidence set size). Lemma D.4 enables us to analyze the leader’s cumulative reward from each arm $a\in\mathcal{A}$ and thus bound the leader’s regret. For the follower’s regret, Lemma D.4 enables us to bound the number of times that arms outside of $\mathcal{A}_{\epsilon}$ are chosen, which enables us to bound the follower’s regret. We defer the full proof to Section D.3. ∎
@@ 4.5 Regret lower bound
[…]
@@
In this sketch, we give intuition for a weaker bound of $\Omega(T^{2/3})$, deferring the strengthening to $\Omega(T^{2/3}\cdot\left(\left|\mathcal{B}\right|\right)^{1/3})$ to Section C.5. Like in the proof of Theorem 3.1, it suffices to consider a centralized environment (Lemma C.1). We show that on the $\mathcal{I}$ and $\tilde{\mathcal{I}}$ in Table 3 (with $\delta=\Theta(T^{-1/3})$), at least one player incurs $\Omega(T^{2/3})$ regret on at least one of these instances. The only way to distinguish the instances is to pull $(a_{1},b_{2})$ at least $\Omega(T^{2/3})$ times, which gives low utility for both players. Intuitively, when the algorithm fails to distinguish $\mathcal{I}$ and $\tilde{\mathcal{I}}$, the algorithm must choose the same distribution over $\mathcal{A}\times\mathcal{B}$, but this gives $\Theta(T^{-1/3})$ loss for the leader on $\mathcal{I}$ or $\Theta(T^{-1/3})$ loss for the follower on $\tilde{\mathcal{I}}$. The full proof, which relies on a KL-divergence argument, is deferred to Section C.5. ∎
@@
5 Relaxed settings that permit faster learning

### v2

@@ 3 Stackelberg value is unachievable
Consider StrongDSGs or WeakDSGs. For any algorithms $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$, there exists an instance $\mathcal{I}^{*}$ with $|\mathcal{A}|=|\mathcal{B}|=2$ where at least one of the players incurs linear regret with respect to the Stackleberg benchmarks $\beta_{1}^{\text{orig}}$ and $\beta_{2}^{\text{orig}}$, that is: $\max(R_{1}(T;\mathcal{I}^{*}),R_{2}(T;\mathcal{I}^{*}))=\Omega(T)$.
@@
It suffices to prove this lower bound in a centralized environment where a single learner can choose action pairs $(a,b)$ and observes rewards for both players (Lemma B.1). We show that on the instances $\mathcal{I}$ and $\tilde{\mathcal{I}}$ in Table 1 (with $\delta=O(1/\sqrt{T})$), at least one of the players incurs linear regret on at least one of the instances. The small value of $\delta$ means that the algorithm fails to distinguish between these instances with constant probability. Nonetheless, the benchmarks are very different: on instance $\mathcal{I}$, $(a^{*},b^{*})=(a_{1},b_{1})$, $\beta_{1}^{\text{orig}}=0.6$ and $\beta_{2}^{\text{orig}}=\delta>0$; on instance $\tilde{\mathcal{I}}$, $(a^{*},b^{*})=(a_{2},b_{1})$, $\beta_{1}^{\text{orig}}=0.5$, and $\beta_{2}^{\text{orig}}=0.6$. Intuitively, when the algorithm fails to distinguish between these instances, then it must choose the same distribution over $\mathcal{A}\times\mathcal{B}$ on both $\mathcal{I}$ and $\tilde{\mathcal{I}}$. However, any such distribution either incurs constant loss for the leader on $\mathcal{I}$ or constant loss for the follower on $\tilde{\mathcal{I}}$. We formalize this proof in Section B.4. ∎
@@ 4.1 $\gamma$-tolerant benchmark
Defining the $\epsilon$-best response set $\mathcal{A}_{\epsilon}$ for the leader is more subtle. Informally, we define this set to include any action $a\in\mathcal{A}$ which has “any chance” of doing at least as well as the the leader’s best action if the follower is $\epsilon$-best responding. Specifically, this includes actions $a$ where some action $b\in\mathcal{B}_{\epsilon}(a)$ achieves utility close to $\max_{a^{\prime}\in\mathcal{A}}\min_{b^{\prime}\in\mathcal{B}_{\epsilon}(a)}v_{1}(a^{\prime},b^{\prime})$ for the leader: [^14] At first glance, it might appear more natural to instead define $\mathcal{A}_{\epsilon}$ to be all actions where the follower’s worst-case approximate best-response yields high utility for the leader, that is: $\{a\in\mathcal{A}\mid\textbf{min}_{b\in\mathcal{B}_{\epsilon}(a)}v_{1}(a,b)\geq\max_{a^{\prime}\in\mathcal{A}}\min_{b^{\prime}\in\mathcal{B}_{\epsilon}(a^{\prime})}v_{1}(a^{\prime},b^{\prime})-\epsilon\}$. However, this set does not necessarily contain the leader’s best-response set $\left\{a\in\text{argmax}_{a^{\prime}\in\mathcal{A}}v_{1}(a^{\prime},b^{*}(a^{\prime}))\right\}$, which makes it a less natural definition of an approximate best-response set.
$\mathcal{A}_{\epsilon}=\{a\in\mathcal{A}\mid\max_{b\in\mathcal{B}_{\epsilon}(a)}v_{1}(a,b)\geq\max_{a^{\prime}\in\mathcal{A}}\min_{b^{\prime}\in\mathcal{B}_{\epsilon}(a^{\prime})}v_{1}(a^{\prime},b^{\prime})-\epsilon\}.$
Observe that the set $\mathcal{B}_{\epsilon}(a)$ always contains the follower’s best-response set $\left\{b\in\text{argmax}_{b^{\prime}\in\mathcal{B}}v_{1}(a,b^{\prime})\right\}$, and furthermore approaches this best-response set in the limit as $\epsilon\rightarrow 0$; similarly, the set $\mathcal{A}_{\epsilon}$ always contains the leader’s best-response set $\left\{a\in\text{argmax}_{a^{\prime}\in\mathcal{A}}v_{1}(a^{\prime},b^{*}(a^{\prime}))\right\}$ where ties are broken in favor of the follower, and furthermore approaches this best-response set in the limit as $\epsilon\rightarrow 0$.
@@ 4.2 Linear regret for ExploreThenCommit
We first show that out-of-box stochastic bandit algorithms do not directly provide sublinear regret against the $\gamma$-tolerant benchmark, where the challenge is that the leader’s learning gets distorted when both players simultaneously learn. To demonstrate this, we consider ExploreThenCommit (Algorithm 1) and show that if both the leader and follower are running this algorithm, the regret could be linear for both players.
@@
             Set $S(c):=\left\{r\mid\exists(t^{\prime},c_{t^{\prime}},r)\in H^{*}\text{ s.t. }c=c_{t^{\prime}}\right\}$
              // observed rewards
             $\hat{v}(c)\leftarrow(\sum_{r\in S(c)}r)/|S(c)|$
              // compute empirical mean
@@
11      return point mass at $\text{argmax}_{c\in\mathcal{C}}\hat{v}(c)$
@@
Consider StrongDSGs where the follower runs a separate instantiation of
@@
The intuition is that in the leader’s exploration phase, the follower alternates uniformly between all actions $\mathcal{B}$. This distorts the leader’s learning during the leader’s exploration phase, and as a result, the leader can choose a highly suboptimal arm during the commit phase. This can lead to linear regret for both players. We construct a single instance (Table 2, with $\delta=0.1$) where both players incur linear regret. The full proof is deferred to Section C.1. ∎
4.3 Warmup Algorithm
The constant regret in Proposition 4.3 motivates the design of more sophisticated algorithms where the leader waits for the follower to partially converge before starting to learn. As a warmup, we show that a simple modification of the setup of Proposition 4.3 guarantees sublinear regret (i.e., $O\left(|\mathcal{A}|^{1/3}|\mathcal{B}|^{1/3}(\log T)^{1/3}T^{2/3}\right)$ regret for both players). In this algorithm, both players run ExploreThenCommit-based algorithms, but the leader waits for the follower to finish exploring before starting to explore. More specifically, the leader runs ExploreThenCommitThrowOut, which acts similar to ExploreThenCommit, but with an extra exploration phase at the start, after which all rewards are thrown out. This initial phase is to allow the follower to partially converge.
@@ 4.3 Warmup Algorithm
       $H^{*}=\left\{(t^{\prime}-E^{\prime}\cdot|\mathcal{C}|,c_{t^{\prime}},r)\mid\exists(t^{\prime},c_{t^{\prime}},r)\in H\text{ s.t. }t^{\prime}>E^{\prime}\cdot|\mathcal{C}|\right\}$
        // Throw out first $E^{\prime}\cdot|\mathcal{C}|$ rounds of history
@@
Consider a StrongDSG where the follower runs a separate instantiation of
@@
The “throw out” phase for $\texttt{ALG}_{1}$ enables $\texttt{ALG}_{2}$ to learn and commit to near-optimal actions. The meaningful exploration for $\texttt{ALG}_{1}$ thus begins after the follower has committed to actions. This enables $\texttt{ALG}_{1}$ to identify a near-optimal action given the arms that $\texttt{ALG}_{2}$ has already committed to after the first phase of exploration. Returning to our $\gamma$-tolerant benchmarks, for each player, we can upper bound regret by setting $\epsilon$ to be the suboptimality of the other player and achieve the desired regret bound. The full proof is deferred to Appendix C.2.
@@
4.4 Main Algorithm
@@
The algorithm $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$ takes as input $E\in[T/\left|\mathcal{A}\right|]$. When $\texttt{ALG}_{1}$ applied to an instance, for the first $|\mathcal{A}|\cdot E$ rounds, the algorithm $\texttt{ALG}_{1}$ pulls each arm in $\mathcal{A}$ a total of $E$ times, and then discards all history from these rounds. For the remaining $T-|\mathcal{A}|\cdot E$ rounds, the algorithm runs UCB, computing the upper confidence bound $v^{\text{UCB}}_{1}(a)=\hat{v}_{1,t}(a)+\alpha_{t}(a)$ using confidence bound $\alpha_{t}(a)=\Theta\left(\sqrt{(\log T)/n_{E\cdot|\mathcal{A}|,t}(a)}\right)$, where $\hat{v}_{1,t}(a)$ is the empirical mean and $n_{E\cdot|\mathcal{A}|,t}(a)$ is the number of times that action $a$ is chosen in the UCB phase after time step $E\cdot|\mathcal{A}|$ and prior to time step $t$. The algorithm then chooses the arm with maximum upper confidence bound.
@@
       $H^{*}=\left\{(t^{\prime}-E\cdot|\mathcal{A}|,a_{t^{\prime}},r)\mid\exists(t^{\prime},a_{t^{\prime}},r)\in H\text{ s.t. }t^{\prime}>E\cdot|\mathcal{A}|\right\}$
        // Throw out first $E\cdot|\mathcal{A}|$ rounds of history
       Initialize $\hat{v}_{1}(a)=1$ for $a\in\mathcal{A}$.
        // Initialize empirical means.
       Initialize $v^{\text{UCB}}_{1}(a)=1$ for $a\in\mathcal{A}$.
        // Initialize UCB.
@@
             Set $S(a):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}}))\in H^{*}\text{ s.t. }a=a_{t^{\prime}},r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}})=r\right\}$
              // Observed rewards
@@
                   $\hat{v}_{1}(a)\leftarrow(\sum_{r\in S(a)}r)/|S(a)|$
                    // Empirical mean
                   $\alpha(a)\leftarrow 10\cdot\frac{\sqrt{\log T}}{\sqrt{|S(a)|}}$
                    // confidence bound width
@@ 10                   $v^{\text{UCB}}_{1}(a)\leftarrow\min(1,\hat{v}_{1}(a)+\alpha(a))$
      Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\left(v^{\text{UCB}}_{1}(a)\right)$.
        // arm with max upper confidence bound
@@
Let $E=\Theta(|\mathcal{A}|^{-2/3}(|\mathcal{B}|\log T)^{1/3}T^{2/3})$. Consider a StrongDSG, where $\texttt{ALG}_{2}$ is any algorithm with high-probability instantaneous regret $g(t,T,\mathcal{B})=O\left((|\mathcal{A}||\mathcal{B}|\log T)^{1/3}T^{-1/3}\right)$ for $t>E$ and $g(t,T,\mathcal{B})=1$ for $t\leq E$, and where $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Then, it holds that the regret with respect to the $\gamma$-tolerant benchmarks $\beta_{1}^{\text{tol}}$ and $\beta_{2}^{\text{tol}}$ is bounded as:
@@
The intuition is that the exploration phase of ExploreThenUCB ensures that all of the follower’s actions have bounded suboptimality, and the UCB phase accounts for the follower changing which action they choose over time. In more detail, high-probability instantaneous regret guarantees that after the explore phase, all actions that the follower’s chooses are within the $\epsilon$-best-response set $\mathcal{B}_{\epsilon^{*}}(a)$ for $\epsilon^{*}=\Theta((|\mathcal{A}|\cdot|\mathcal{B}|\cdot\log T)^{1/3}T^{-1/3})$. For the UCB phase, the main lemma (Lemma C.4) is that if an arm $a\in\mathcal{A}$ is pulled, the empirical mean is at least $\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\Theta\left(\sqrt{\log T/n_{E\cdot|\mathcal{A}|,t}(a)}\right)$ (the optimal utility for the leader when the follower worst-case $\epsilon$-best-responds minus the confidence set size). Lemma C.4 enables us to analyze the leader’s cumulative reward from each arm $a\in\mathcal{A}$ and thus bound the leader’s regret. For the follower’s regret, Lemma C.4 enables us to bound the number of times that arms outside of $\mathcal{A}_{\epsilon}$ are chosen, which enables us to bound the follower’s regret. We defer the full proof to Section C.3. ∎
@@ 4.5 Regret lower bound
Consider StrongDSGs or WeakDSGs with action sets $\mathcal{A}$ and $\mathcal{B}$ such that $|\mathcal{A}|\geq 2$ and $|\mathcal{B}|\geq 2$. For any algorithms $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$, there exists an instance $\mathcal{I}^{*}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ such that at least one of the players incurs $\Omega(T^{2/3}\cdot\left(\left|\mathcal{B}\right|\right)^{1/3})$ regret with respect to the $\gamma$-tolerant benchmarks $\beta_{1}^{\text{tol}}$ and $\beta_{2}^{\text{tol}}$:
@@
In this sketch, we give intuition for a weaker bound of $\Omega(T^{2/3})$, deferring the strengthening to $\Omega(T^{2/3}\cdot\left(\left|\mathcal{B}\right|\right)^{1/3})$ to Section B.5. Like in the proof of Theorem 3.1, it suffices to consider a centralized environment (Lemma B.1). We show that on the $\mathcal{I}$ and $\tilde{\mathcal{I}}$ in Table 3 (with $\delta=\Theta(T^{-1/3})$), at least one player incurs $\Omega(T^{2/3})$ regret on at least one of these instances. The only way to distinguish the instances is to pull $(a_{1},b_{2})$ at least $\Omega(T^{2/3})$ times, which gives low utility for both players. Intuitively, when the algorithm fails to distinguish $\mathcal{I}$ and $\tilde{\mathcal{I}}$, the algorithm must choose the same distribution over $\mathcal{A}\times\mathcal{B}$, but this gives $\Theta(T^{-1/3})$ loss for the leader on $\mathcal{I}$ or $\Theta(T^{-1/3})$ loss for the follower on $\tilde{\mathcal{I}}$. The full proof, which relies on a KL-divergence argument, is deferred to Section B.5. ∎
@@
5 Relaxed Settings with Faster Learning